# The stringy geometry of integral cohomology in mirror symmetry

### v2

@@ 1 Introduction
At the heart of this remarkable duality is closed string mirror symmetry: an isomorphism of two superconformal field theories: the (2,2) superconformal field theory (SCFT) ${\cal C}[X]$ based on a non-linear sigma model (NLSM) with target space $X$ and another (2,2) SCFT ${\cal C}[X^{\circ}]$ based on a NLSM with target space $X^{\circ}$. The exchange of the Hodge numbers is the most basic and familiar aspect of mirror symmetry. It has an immediate and satisfying reflection in the SCFT: $h^{1,2}(X)$ encodes the number of exactly marginal operators with charges $q_{\text{\tiny{L}}}=q_{\text{\tiny{R}}}=1$ with respect to the $\operatorname{U{}}(1)_{\text{L}}\times\operatorname{U{}}(1)_{\text{R}}$ R-symmetry of the SCFT (these are the “cc deformations”), while $h^{1,1}(X)$ encodes the exactly marginal operators with charges $-q_{\text{\tiny{L}}}=q_{\text{\tiny{R}}}=1$ (these are the “ac deformations”). Integrating these deformations we obtain the conformal manifold ${\cal M}(X)={\cal M}_{\text{cc}}(X)\times{\cal M}_{\text{ac}}(X)$ of the SCFT, where each factor is a special Kähler manifold with respect to the Zamolodchikov metric.
@@
$A(X)$ =$=\{H^{2}(X,{\mathbb{Z}})\}_{\text{tor}}\quad\text{and}\quad B(X)=\{H^{3}(X,{\mathbb{Z}})\}_{\text{tor}}~{}.$ (1.1)
@@
00{0}${H^{2}(X,{\mathbb{R}})/H^{2}(X,{\mathbb{Z}})}$${H^{2}(X,\operatorname{U{}}(1))}$${B(X)}$${0~{}.}$ (1.3)
The middle term has the intuitive interpretation as a choice of phases assigned by the NLSM path integral to topological sectors labeled by the homology classes $[f(\Sigma)]\in H_{2}(X,{\mathbb{Z}})$, where $f:\Sigma\to X$ is the map from the worldsheet $\Sigma$ to the target space $X$. Thus $B(X)$ describes the equivalence classes of flat gerbes with respect to shifts of $B$—the connection on the gerbe—by closed $2$-forms.
@@ 2 A review of topology and geometry
Let $X$ be a Calabi-Yau manifold of complex dimension $d$ [^2] There are a few slightly different definitions of Calabi-Yau manifolds, and we choose one that is general enough to cover the examples of interest to us, while excluding special cases like $T^{6}$. Our presentation of the geometry mostly follows [30] and [31], to which we refer the reader for further discussion and references. That is, $X$ is a compact Kähler manifold with $H^{i}(X,{\cal O}_{X})=0$ for $0<i<d$, while $H^{0}(X,{\cal O}_{X})$ and $H^{d}(X,{\cal O}_{X})$ are each isomorphic to ${\mathbb{C}}$. It follows from Yau’s theorem that $X$ admits a Ricci-flat metric $g$ with holonomy group $\text{Hol}(g)\subseteq\operatorname{SU}(d)$.
@@ 2.1 Holonomy and the fundamental group
    with $\zeta_{k}=e^{2\pi i/k}$ a primitive $k$-th root of unity. The action has $5$ fixed points in ${\mathbb{P}}^{4}$, where all but one of the projective coordinates $Z_{i}$ is zero, but these are missed by a generic $G$–invariant hypersurface. [^4] One does not have to look far for a representative smooth hypersurface: the Fermat quintic does the job. Since the $G$-action on $X$ is free, we have $\pi_{1}(X)=G={\mathbb{Z}}_{5}$.
@@
    from which we also read off $h^{1,1}=1$ and $h^{1,2}=21$; the former is the Kähler deformation inherited from the quintic, and the latter are $21$ complex structure deformations, each of which can be represented by a deformation of the $G$–invariant polynomial. With our definition of the independent torsion subgroups we have $A(X)={\mathbb{Z}}_{5}$, and $B(X)=0$. We will return to this example repeatedly in what follows.
@@ 2.3 Torsion in K-theory
This is not yet enough to demonstrate that the sequence for $\{K^{0}(X)\}_{\text{tor}}$ splits. The desired splitting is equivalent to the existence of a group homomorphism $h:A(X)\to\{K^{0}(X)\}_{\text{tor}}$ such that $c_{1}(h(\omega))=\omega$ for all $\omega\in A(X)$. We have not succeeded in showing directly that $h$ exists, as it seems to depend on a certain quadratic refinement in torsional cohomology. [^8] This quadratic refinement holds if $c_{1}({\cal L})\cup c_{1}({\cal L})=0\mod 2$ for all torsion line bundles. We checked that this condition is satisfied for the $X_{2}$ and $X_{3}$ examples of section 2.1. Instead, we will see that its existence follows from the properties of $K^{1}(X)$ and a perfect pairing between the torsion K-groups.
Observations on $K^{1}(X)$
@@
00{0}${H^{5}(X,{\mathbb{Z}})}$${K^{1}(X)}$${H^{3}(X,{\mathbb{Z}})}$${0~{}.}$ (2.26)
Since $H^{k}(X,{\mathbb{Z}})\simeq H^{k+1}(\Sigma X,{\mathbb{Z}})$ for $k>0$, we obtain
00{0}${H^{6}(\Sigma X,{\mathbb{Z}})}$${K^{1}(X)}$${H^{4}(\Sigma X,{\mathbb{Z}})}$${0~{}.}$$\scriptstyle{c_{2}}$ (2.27)
We want to argue that for every $\upsilon\in H^{4}(\Sigma X,{\mathbb{Z}})$ there exists a virtual bundle ${\cal E}_{\upsilon}\in\widetilde{K}^{0}(\Sigma X)\simeq K^{1}(X)$ such that $c_{2}({\cal E}_{\upsilon})=\upsilon$, $c_{1}({\cal E}_{\upsilon})=0$, and $c_{3}({\cal E}_{\upsilon})=0$. We cannot just apply the previous argument to reach this conclusion, because $\Sigma X$ need not be a manifold, and therefore the concept of spin need not be well-defined. Instead we consider the following diagram:
[…]
where $S$ is the isomorphism induced by the suspension. The desired result now follows from two observations: (i) since $X$ is spin $\text{Sq}^{2}=0$ on $H^{3}(X,{\mathbb{Z}})$; (ii) the Steenrod square operation is stable with respect to suspension [47].
Let $h^{\prime}:H^{4}(\Sigma X,{\mathbb{Z}})\to K^{1}(X)$ be the map $h^{\prime}(\upsilon)={\cal E}_{\upsilon}$. Then
$h^{\prime}(\upsilon_{1}+\upsilon_{2})$ $={\cal E}_{\upsilon_{1}+\upsilon_{2}}~{},$ $h^{\prime}(\upsilon_{1})\oplus h^{\prime}(\upsilon_{2})={\cal E}_{\upsilon_{1}}\oplus{\cal E}_{\upsilon_{2}}~{}.$ (2.29)
It follows that
$c({\cal E}_{\upsilon_{1}+\upsilon_{2}})=1+\upsilon_{1}+\upsilon_{2}=c({\cal E}_{\upsilon_{1}}\oplus{\cal E}_{\upsilon_{2}})~{}.$ (2.30)
Although this shows that the Chern classes of ${\cal E}_{\upsilon_{1}+\upsilon_{2}}$ agree with those of ${\cal E}_{\upsilon_{1}}\oplus{\cal E}_{\upsilon_{2}}$, the two need not be isomorphic in $\widetilde{K}^{0}(\Sigma X)=K^{1}(X)$ because there may exist stable non-trivial vectors bundles that are not detected by characteristic classes. For instance, there is a non-trivial torsion vector bundle on ${\mathbb{R}}{\mathbb{P}}^{7}$ with all vanishing characteristic classes; this example is discussed on page 10 of [48].
But now, suppose that every element $x\in A(X)^{\ast}$ is of odd order. Then its mod $2$ reduction as characteristic class is necessarily trivial. Thus, every $x\in A(X)^{\ast}$ can be realized as $c_{3}({\cal E})$ of a virtual vector bundle ${\cal E}\in\widetilde{K^{0}}(\Sigma X)$, and therefore (2.27) implies that ${\cal E}$ is trivial if and only if all of its Chern classes vanish.
It follows that $h^{\prime}(\upsilon_{1})\oplus h^{\prime}(\upsilon_{2})$ and $h^{\prime}(\upsilon_{1}+\upsilon_{2})$ are equivalent in $K^{1}(X)$. Thus, $h^{\prime}$ is a group homomorphism, and because $c_{2}(h^{\prime}(\upsilon))=\upsilon$, it gives the desired splitting of the sequence. Taking the torsion subgroups, we therefore have
@@
Returning to our discussion of $\{K^{0}(X)_{\text{tor}}$, we recall that there is a perfect pairing (see section 3.1 of  [48]):
$\{K^{0}(X)\}_{\text{tor}}\times\{K^{1}(X)\}_{\text{tor}}\to\operatorname{U{}}(1)~{}.$ (2.32)
This means $\{K^{0}(X)\}_{\text{tor}}$ is the Pontryagin dual to $K^{1}(X)$, and therefore
$\{K^{0}(X)\}_{\text{tor}}=A(X)\oplus B(X)^{\ast}~{}.$ (2.33)
@@ 3.1 Intuition from the non-linear sigma model
×$\qquad\qquad\qquad\qquad\qquad\times\exp\left[2\pi{\mathrm{i}}\left(\int_{C_{\text{f}}}f^{\ast}(\beta_{\text{f}})+\langle C_{\text{t}},\beta_{\text{t}}\rangle\right)\right]~{}.$ (3.1)
@@ 3.3 Mirror symmetry
The linear sigma model is obtained as follows. There are $n$ (2,2) chiral superfields $Z_{\rho}$, and these are minimally coupled to $n-d$ abelian vector multiplets, with charges determined by ${\cal G}$. We introduce an additional chiral superfield $\Phi_{0}$ with gauge charges chosen such that $\Phi_{0}\prod_{\rho}Z_{\rho}$ is gauge-invariant, and we set the chiral superpotential to be
@@
The $\alpha_{m}$ are generic complex coefficients, and $\langle m,\rho\rangle$ denotes the pairing between the $M$ and $N$ lattices. The remaining term in the Lagrangian is the twisted chiral superpotential, which is linear in the gauge-field strengths of the vector multiplets and introduces $n-d$ Fayet-Iliopoulos parameters $r_{\alpha}$ complexified by $n-d$ $\theta$-angles $\theta_{\alpha}$. The theory has a non-anomalous $\operatorname{U{}}(1)_{\text{L}}\times\operatorname{U{}}(1)_{\text{R}}$ R-symmetry, and for a suitable choice of the $r_{\alpha}$, it is believed to flow at low energy to the same fixed point as the NLSM based on the CY hypersurface $\overline{X}$ defined by the locus
@@
Having glanced at the essential features of the Batyrev mirror symmetry, we proceed to study our example. Starting with the mirror pair $\overline{X}$, $\overline{X}^{\circ}$, we observe that we can tune the coefficients $\alpha_{m}$ for the quintic polynomial in ${\mathbb{P}}^{4}$ so that the linear sigma model (and the geometry) acquires an additional $G={\mathbb{Z}}_{5}$ symmetry, which acts by phases as in (2.1). The action on a monomial corresponding to lattice point $m\in\Delta$ is to multiply it by a phase $\zeta_{5}^{\langle m,\rho_{g}\rangle}$, where $\rho_{g}=\rho_{1}+2\rho_{2}+3\rho_{3}+4\rho_{4}$. We say $m$ is a $G$–invariant lattice point if $\langle m,\rho_{g}\rangle=0\mod 5$, and it is not hard to see that there are $26$ of these. They consist of the origin, the vertices $v_{0},\ldots,v_{4}$, as well points contained in the relative interior of dimension $2$ faces $F_{abc}$
@@
Once we set $\alpha_{m}=0$ for all non-invariant lattice points and keep the remaining $\alpha_{m}$ generic, we obtain a family of quintic hypersurfaces $\overline{X}\subset{\mathbb{P}}^{4}$, each with a freely-acting isometry $G$. The moduli space of these theories has the form ${\cal M}_{\text{ac}}(\overline{X})\times{\cal M}_{\text{cc}}(\overline{X})_{G}$, where the subscript indicates that we restrict to the $G$–invariant locus in the quintic moduli space. Note that this places no restriction on the complexified Kähler parameter. Taking the quotient, we obtain our example $X$ with $h^{1,1}(X)=1$, $h^{1,2}(X)=21$, $A(X)=G^{\ast}$, and $B(X)=0$. By the arguments in the previous section the SCFT moduli space is given by ${\cal M}_{\text{ac}}(X)\times{\cal M}_{\text{cc}}(X)={\cal M}_{\text{ac}}(\overline{X})\times{\cal M}_{\text{cc}}(\overline{X})_{G}$.
@@
The mirror of $A(X)$—general story
Our example is an explicit realization of a general phenomenon, where mirror symmetry can be used to relate the torsion group $A(X)$ to a global symmetry in the mirror description. Suppose $X$ and $X^{\circ}$ are a mirror pair of smooth Calabi-Yau $3$-folds, and $\pi_{1}(X)=G$ is a finite abelian group. $X$ has a universal cover $\overline{X}$ with a group $G$ of Kähler isometries such that ${\cal C}[\overline{X}]$ has $G$ as a global symmetry group on an appropriate locus in the SCFT moduli space. We denote this restriction by the notation ${\cal C}[\overline{X}_{G}]$. If we suppose further that $\overline{X}$ has a known mirror $\overline{X}^{\circ}$, we can use the mirror isomorphism to obtain a family of mirror theories ${\cal C}[(\overline{X}_{G})^{\circ}]$ with a global symmetry $G^{\circ}\simeq G$ [^20] Our abstruse notation ${\cal C}[(\overline{X}_{G})^{\circ}]$ denotes the family of SCFTs mirror to the family ${\cal C}[\overline{X}_{G}]$. As the example already shows, it need not correspond to any smooth geometry, and, indeed, it need not have a geometric interpretation at all, since the $G$—invariant locus may not contain a large complex structure point. We then obtain the following diagram relating various SCFTs:
@@
$G^{\circ}$ =$=\{\overline{\mu}g\overline{\mu}^{-1}~{}~{}~{}|~{}~{}g\in G\}~{},$ $(G_{\text{q}})^{\circ}$ =$=\{\mu g\mu^{-1}~{}~{}~{}|~{}~{}g\in G_{\text{q}}\}~{}.$ (3.24)
In our explicit example we found that the mirror dual of $A(X)$—a non-geometric symmetry of ${\cal C}[X]$—is a geometric symmetry of ${\cal C}[X^{\circ}]$. The general structure and the special features of the example lead to a number of additional questions. For instance, the states charged with respect to $A(X)$ are string winding modes, and intuition from T-duality would suggest that their duals are some “momentum” states that could perhaps be represented by eigenmodes of an appropriate Laplace operator defined on $X^{\circ}$ and graded by representations of the geometric symmetry $(G_{\text{q}})^{\circ}$. It would be useful to make this correspondence precise; perhaps it can be explored using recent advances in numeric methods for CY geometry such as [76, 77]. Further work is also needed to study the correspondence when $\pi_{1}(X)$ is non-abelian, where more sophisticated categorical methods would need to be used to recover the “upstairs” SCFTs from ${\cal C}[X]$ and ${\cal C}[X^{\circ}]$.

### v3

@@ 1 Introduction
At the heart of this remarkable duality is closed string mirror symmetry: an isomorphism of two superconformal field theories: the (2,2) superconformal field theory (SCFT) ${\cal C}[X]$ based on a non-linear sigma model (NLSM) with target space $X$ and another (2,2) SCFT ${\cal C}[X^{\circ}]$ based on a NLSM with target space $X^{\circ}$. The exchange of the Hodge numbers is the most basic and familiar aspect of mirror symmetry. It has an immediate and satisfying reflection in the SCFT: $h^{1,2}(X)$ encodes the number of exactly marginal operators with charges $q_{\text{\tiny{L}}}=q_{\text{\tiny{R}}}=1$ with respect to the $\operatorname{U{}}(1)_{\text{L}}\times\operatorname{U{}}(1)_{\text{R}}$ R-symmetry of the SCFT (these are the “cc deformations”), while $h^{1,1}(X)$ encodes the exactly marginal operators with charges $q_{\text{\tiny{L}}}=q_{\text{\tiny{R}}}=1$ (these are the “ac deformations”). Integrating these deformations we obtain the conformal manifold ${\cal M}(X)={\cal M}_{\text{cc}}(X)\times{\cal M}_{\text{ac}}(X)$ of the SCFT, where each factor is a special Kähler manifold with respect to the Zamolodchikov metric.
@@
$A(X)$ $=\{H^{2}(X,{\mathbb{Z}})\}_{\text{tor}}\quad\text{and}\quad B(X)=\{H^{3}(X,{\mathbb{Z}})\}_{\text{tor}}~.$ (1.1)
@@
Returning to the closed string sector, the significance of $A(X)$ is fairly clear from its interpretation as the Pontryagin dual of the abelianization of the fundamental group $\pi_{1}(X)$. Since we can construct $X$ as a quotient $\overline{X}/\Gamma$, where $\overline{X}$ is an isometric universal cover of $X$ that admits the free action of a group $\Gamma\simeq\pi_{1}(X)$, ${\cal C}[X]$ is an orbifold conformal field theory: ${\cal C}[X]={\cal C}[\overline{X}]/\Gamma$. If $X$ has a known mirror $X^{\circ}$ we can find a theory $\overline{{\cal C}}^{\circ}$ with a global symmetry $\Gamma^{\circ}$ isomorphic to $\Gamma$ such that $\overline{{\cal C}}^{\circ}/\Gamma^{\circ}={\cal C}[X^{\circ}]$. When $\Gamma$ is finite and abelian, this follows from standard properties of the quantum symmetry associated to an abelian orbifold, but for more general finite groups it requires a categorical symmetry construction—reviewed and axiomatized in [10, 11]. [^1] If $\Gamma$ is of infinite order, then $X$ is a quotient of $T^{6}$ or $\text{K3}\times T^{2}$ by a freely–acting group, and the possibilities are classified in [12]. These theories should be considered in its own right for many purposes, including mirror symmetry, which is now inherited from simpler operations in the cover [13]. In either case, however, there is no guarantee that $\overline{{\cal C}}^{\circ}$ can be presented as ${\cal C}[\overline{Y}]$ for some CY $\overline{Y}$; as we show, even when it can be so presented, the action of the group $\Gamma^{\circ}$ may not have a geometric interpretation.
@@
The middle term has the intuitive interpretation as a choice of phases assigned by the NLSM path integral to topological sectors labeled by the homology classes $f(\Sigma)]\in H_{2}(X,{\mathbb{Z}})$, where $f:\Sigma\to X$ is the map from the worldsheet $\Sigma$ to the target space $X$. Thus $B(X)$ describes the equivalence classes of flat gerbes with respect to shifts of BB—the connection on the gerbe—by closed $2$-forms.
@@ 2 A review of topology and geometry
Let $X$ be a Calabi-Yau manifold of complex dimension dd. [^2] There are a few slightly different definitions of Calabi-Yau manifolds, and we choose one that is general enough to cover the examples of interest to us, while excluding special cases like $T^{6}$. Our presentation of the geometry mostly follows [30] and [31], to which we refer the reader for further discussion and references. That is, $X$ is a compact Kähler manifold with $H^{i}(X,{\cal O}_{X})=0$ for $0<i<d$, while $H^{0}(X,{\cal O}_{X})$ and $H^{d}(X,{\cal O}_{X})$ are each isomorphic to $\mathbb{C}$}. It follows from Yau’s theorem that $X$ admits a Ricci-flat metric $g$ with holonomy group $\text{Hol}(g)\subseteq\operatorname{SU}(d)$.
@@ 2.1 Holonomy and the fundamental group
    with $\zeta_{k}=e^{2\pi i/k}$ a primitive $k$-th root of unity. The action has $5$ fixed points in ${\mathbb{P}}^{4}$, where all but one of the projective coordinates $Z_{i}$ is zero, but these are missed by a generic GG–invariant hypersurface. [^4] One does not have to look far for a representative smooth hypersurface: the Fermat quintic does the job. Since the $G$-action on $X$ is free, we have $\pi_{1}(X)=G={\mathbb{Z}}_{5}$.
@@
    from which we also read off $h^{1,1}=1$ and $h^{1,2}=21$; the former is the Kähler deformation inherited from the quintic, and the latter are $21$ complex structure deformations, each of which can be represented by a deformation of the GG–invariant polynomial. With our definition of the independent torsion subgroups we have $A(X)={\mathbb{Z}}_{5}$, and $B(X)=0$. We will return to this example repeatedly in what follows.
@@ 2.3 Torsion in K-theory
Splitting the sequences
The preceding observations do not by imply that the sequence for $\{K^{0}(X)\}_{\text{tor}}$ splits. The desired splitting is equivalent to the existence of a surjective group homomorphism $h:\{K^{0}(X)\}_{\text{tor}}\to H^{4}(X,{\mathbb{Z}})$. Taking rational coefficients the Chern character provides the group isomorphism $\operatorname{ch}:K^{0}(X)\otimes_{{\mathbb{Z}}}{\mathbb{Q}}\to H^{\text{even}}(X,{\mathbb{Q}})$ [46]. Restricting this to the second Chern character we find the surjective map $\operatorname{ch}_{2}:K^{0}(X)\otimes_{{\mathbb{Z}}}{\mathbb{Q}}\to H^{4}(X,{\mathbb{Q}})$, with image
$\operatorname{ch}_{2}(E)={\textstyle\frac{1}{2}}c_{1}(E)\cup c_{1}(E)-c_{2}(E)~.$ (2.26)
But now we observe that if we restrict to torsion subgroups, then our assumption that $H^{\ast}(X,{\mathbb{Z}})_{\text{tor}}$ contains no generators of even order implies that for every $E\in\{K^{0}(X)\}_{\text{tor}}$ $c_{1}(E)\cup c_{1}(E)$ is an even class in $B(X)^{\ast}$, and that means there is a surjective group homomorphism
$g$ $:K^{0}(X)_{\text{tor}}\to B(X)^{\ast}~,$ $g(E)={\textstyle\frac{1}{2}}c_{1}(E)\cup c_{1}(E)-c_{2}(E)~.$ (2.27)
Moreover, by observation 3 above this map remains surjective even when restricted to $\operatorname{im}f=\ker c_{1}$, and that means $\phi=g\circ f$ is a surjective finite group endomorphism $\phi:B(X)^{\ast}\to B(X)^{\ast}$ and hence an isomorphism, so that we obtain the split sequence
@@
The sequence for $K^{1}(X)_{\text{tor}}$ splits as well. This follow from the perfect pairing (see section 3.1 of  [48]):
$\{K^{0}(X)\}_{\text{tor}}\times\{K^{1}(X)\}_{\text{tor}}\to\operatorname{U{}}(1)~,$ (2.29)
which implies $\{K^{0}(X)\}_{\text{tor}}$ is the Pontryagin dual to $K^{1}(X)$, and therefore
@@
Having shown the split, we note that our assumptions and mirror symmetry imply that $A(X^{\circ})^{\ast}\subset K^{1}(X^{\circ})_{\text{tor}}$ and $B(X^{\circ})^{\ast}\subset K^{0}(X^{\circ})_{\text{tor}}$ also only have elements of odd degree, and thus the sequences mirror to (2.21) split as well. Thus, mirror symmetry and absence of even order torsion cocycles in $H^{\ast}(X,{\mathbb{Z}})$ imply (1.2).
@@ 3.1 Intuition from the non-linear sigma model
$\qquad\qquad\qquad\qquad\qquad\times\exp\left[2\pi{\mathrm{i}}\left(\int_{C_{\text{f}}}\beta_{\text{f}}+\langle C_{\text{t}},\beta_{\text{t}}\rangle\right)\right]~.$ (3.1)
@@ 3.3 Mirror symmetry
The linear sigma model is obtained as follows. There are $n$ (2,2) chiral superfields $Z_{\rho}$, and these are minimally coupled to ndn-d abelian vector multiplets, with charges determined by $\cal G$}. We introduce an additional chiral superfield $\Phi_{0}$ with gauge charges chosen such that $\Phi_{0}\prod_{\rho}Z_{\rho}$ is gauge-invariant, and we set the chiral superpotential to be
@@
The $\alpha_{m}$ are generic complex coefficients, and $\langle m,\rho\rangle$ denotes the pairing between the $M$ and $N$ lattices. The remaining term in the Lagrangian is the twisted chiral superpotential, which is linear in the gauge-field strengths of the vector multiplets and introduces ndn-d Fayet-Iliopoulos parameters $r_{\alpha}$ complexified by ndn-d $\theta$-angles $\theta_{\alpha}$. The theory has a non-anomalous $\operatorname{U{}}(1)_{\text{L}}\times\operatorname{U{}}(1)_{\text{R}}$ R-symmetry, and for a suitable choice of the $r_{\alpha}$, it is believed to flow at low energy to the same fixed point as the NLSM based on the CY hypersurface $\overline{X}$ defined by the locus
@@
Having glanced at the essential features of the Batyrev mirror symmetry, we proceed to study our example. Starting with the mirror pair $\overline{X}$, $\overline{X}^{\circ}$, we observe that we can tune the coefficients $\alpha_{m}$ for the quintic polynomial in ${\mathbb{P}}^{4}$ so that the linear sigma model (and the geometry) acquires an additional $G={\mathbb{Z}}_{5}$ symmetry, which acts by phases as in (2.1). The action on a monomial corresponding to lattice point $m\in\Delta$ is to multiply it by a phase $\zeta_{5}^{\langle m,\rho_{g}\rangle}$, where $\rho_{g}=\rho_{1}+2\rho_{2}+3\rho_{3}+4\rho_{4}$. We say $m$ is a GG–invariant lattice point if $\langle m,\rho_{g}\rangle=0\mod 5$, and it is not hard to see that there are $26$ of these. They consist of the origin, the vertices $v_{0},\ldots,v_{4}$, as well points contained in the relative interior of dimension $2$ faces $F_{abc}$
@@
Once we set $\alpha_{m}=0$ for all non-invariant lattice points and keep the remaining $\alpha_{m}$ generic, we obtain a family of quintic hypersurfaces $\overline{X}\subset{\mathbb{P}}^{4}$, each with a freely-acting isometry $G$. The moduli space of these theories has the form ${\cal M}_{\text{ac}}(\overline{X})\times{\cal M}_{\text{cc}}(\overline{X})_{G}$, where the subscript indicates that we restrict to the GG–invariant locus in the quintic moduli space. Note that this places no restriction on the complexified Kähler parameter. Taking the quotient, we obtain our example $X$ with $h^{1,1}(X)=1$, $h^{1,2}(X)=21$, $A(X)=G^{\ast}$, and $B(X)=0$. By the arguments in the previous section the SCFT moduli space is given by ${\cal M}_{\text{ac}}(X)\times{\cal M}_{\text{cc}}(X)={\cal M}_{\text{ac}}(\overline{X})\times{\cal M}_{\text{cc}}(\overline{X})_{G}$.
@@
The mirror of A(X)A(X)—general story
Our example is an explicit realization of a general phenomenon, where mirror symmetry can be used to relate the torsion group $A(X)$ to a global symmetry in the mirror description. Suppose $X$ and $X^{\circ}$ are a mirror pair of smooth Calabi-Yau $3$-folds, and $\pi_{1}(X)=G$ is a finite abelian group. $X$ has a universal cover $\overline{X}$ with a group $G$ of Kähler isometries such that ${\cal C}[\overline{X}]$ has $G$ as a global symmetry group on an appropriate locus in the SCFT moduli space. We denote this restriction by the notation ${\cal C}[\overline{X}_{G}]$. If we suppose further that $\overline{X}$ has a known mirror $\overline{X}^{\circ}$, we can use the mirror isomorphism to obtain a family of mirror theories ${\cal C}[(\overline{X}_{G})^{\circ}]$ with a global symmetry $G^{\circ}\simeq G$. [^16] Our abstruse notation ${\cal C}[(\overline{X}_{G})^{\circ}]$ denotes the family of SCFTs mirror to the family ${\cal C}[\overline{X}_{G}]$. As the example already shows, it need not correspond to any smooth geometry, and, indeed, it need not have a geometric interpretation at all, since the GG—invariant locus may not contain a large complex structure point. We then obtain the following diagram relating various SCFTs:
@@
$G^{\circ}$ $=\{\overline{\mu}g\overline{\mu}^{-1}~~~|~~g\in G\}~,$ $(G_{\text{q}})^{\circ}$ $=\{\mu g\mu^{-1}~~~|~~g\in G_{\text{q}}\}~.$ (3.24)
In our explicit example we found that the mirror dual of A(X)A(X)—a non-geometric symmetry of ${\cal C}[X]$—is a geometric symmetry of ${\cal C}[X^{\circ}]$. The general structure and the special features of the example lead to a number of additional questions. For instance, the states charged with respect to $A(X)$ are string winding modes, and intuition from T-duality would suggest that their duals are some “momentum” states that could perhaps be represented by eigenmodes of an appropriate Laplace operator defined on $X^{\circ}$ and graded by representations of the geometric symmetry $(G_{\text{q}})^{\circ}$. It would be useful to make this correspondence precise; perhaps it can be explored using recent advances in numeric methods for CY geometry such as [76, 77]. Further work is also needed to study the correspondence when $\pi_{1}(X)$ is non-abelian, where more sophisticated categorical methods would need to be used to recover the “upstairs” SCFTs from ${\cal C}[X]$ and ${\cal C}[X^{\circ}]$.